# Measuring Sound Symbolism in Audio-Visual Models

### v1

@@ 1 Introduction
Linguistic theory has long posited that the relationship between sound and meaning is arbitrary. This notion, originating from Hermogenes’ discussion with Socrates and later revitalized by John Locke [1], has profoundly influenced modern linguistics. Central to this perspective is Ferdinand de Saussure, widely considered “the father of modern linguistics.” Saussure stated that“le signe est arbitraire111“the sign is arbitrary”,” arguing that the association between words and their referents is based solely on social consensus. He further contended that the meaning of words arises only through their distinctions from one another—hence, a word like “dog” signifies its object only by not being “cat” or “horse.”
Despite the dominance of this view, considerable efforts in recent decades have documented numerous instances challenging the arbitrariness of linguistic signs. Known as sound symbolism, this line of research suggests a systematic and non-arbitrary association between speech sounds and their meanings. Profound cases of sound symbolism include onomatopoeia [2], where the words themselves, such as “quack” and “boom”, mimic actual sounds. Additionally, phonaesthemes [3] suggest that sub-morphemic sequences of sounds convey meanings. For example, in English, the prefix “gl-” appears in several words related to light or vision, such as “gleam,” “glow,” and “glare.” Among the most well-known and thoroughly studied cases is the kiki-bouba effect [4, 5], in which the sharp-sounding “kiki” is typically associated with spiky shapes and the softer “bouba” with rounded forms (as shown in Figure 1). This effect, which demonstrates a non-arbitrary link between speech sounds and visual objects, has proven robust across various languages and cultures [4, 6, 7].
In the field of deep learning, there has been significant interest in exploring the interplay between auditory and visual perceptions, two critical modalities in human sensory experience. This research area, known as audio-visual learning [8], seeks to overcome the limitations of learning from a single modality, thus opening up various new research avenues. A primary focus within this domain is audio-visual representation learning, which aims to develop joint audio and visual representations. Techniques in this field typically involve pretraining models to maximize the mutual information between auditory and visual inputs (whether images or videos). By training with diverse audio-visual datasets, different models achieve various specializations, such as speech-image retrieval [9, 10], visual speech recognition [11], audio event understanding [12, 13]. These pre-trained models can extract highly informative, modality-agnostic representations from raw data, making them beneficial for solving audio-visual tasks even with limited labeled data.
@@
Given their success, researchers are increasingly interested in probing the emergent properties of these pre-trained audio-visual models, such as their potential for acoustic unit discovery [11], word discovery [14], audio-visual localization [12], and video reconstruction [13], in contrast to single-modality models. These interesting capabilities raise a pertinent question: Can these models also capture non-arbitrary connections between sounds and meanings, similar to humans? In other words, does sound symbolism also exist in these pre-trained audio-visual models? In this work, we specifically examine whether pre-trained audio-visual models exhibit the kiki-bouba effect by analyzing how embeddings align across different modalities. Although this iconicity is not essential for the model’s functionality, it could provide insights into the cognitive architectures and learning strategies shared by humans and machines. This would help us understand whether these models interpret novel language in ways consistent with human comprehension, which is the ultimate goal of language understanding. Our findings, regardless of outcome, provide a novel computational perspective on sound symbolism, potentially challenging or reinforcing previous assumptions in the field.
We start with curating a specialized dataset that includes synthesized images and audio samples, each distinctly categorized as ”sharp” or ”round.” We then use a non-parametric approach in a zero-shot setting to probe the inherent knowledge of these models. The experiment results reveal a significant correlation between the models’ outputs and established patterns of the kiki-bouba effect in some audio-visual models. Specifically, audio-visual models trained on spoken image captions can group visual stimuli and audio of different appearances significantly better than chance. Additionally, audio-visual models generally show a more pronounced sound symbolism effect than their purely textual vision-language model counterpart. Our findings shows synergy with existing psychological literature [15, 16, 5, 17, 18] and further support the non-arbitrariness of human language.
@@ 2.3 Interpreting Deep Learning Models
Deep learning has undoubtedly achieved remarkable success across various fields [35, 36, 37]. Researchers are increasingly interested in interpreting these models to better understand the underlying mechanics and rationales behind their decisions [38]. This interpretation often involves feature visualization [39, 40, 41], saliency methods [42, 43, 41], and probing tasks [44, 45, 46, 47, 48]. Recently, the natural language processing community has shown a growing interest in comparing the behavior of vision-language models to aspects of human language processing through probing tasks. These tasks assess models on compositionality [49], abstraction [50], and sound symbolism [51].
In this work, we specifically aim to study the effect of sound symbolism in audio-visual models. By integrating insights from the interpretability of deep learning models, we explore how these black-box systems may encode and reflect non-arbitrary associations between sounds and visual representations akin to humans.
@@ 3 Dataset Collection
To assess the presence of sound symbolism in audio-visual models, we need to inspect how models behave with respect to sound and visual stimuli categorized into sharp or round groups. Therefore, we first construct a specialized dataset with images and audios grounded and grouped by human preference. While it is feasible to use controlled examples from original kiki-bouba experiments (as illustrated in Figure 1), the limited diversity in these samples could potentially undermine the reliability of our evaluation results. Hence, we enhanced the diversity by synthesizing data with generative models, as detailed in Sections 3.1 and 3.2. Overall, our dataset consists of 500 images and 3888 audio samples, each categorized into sharp (denoted as $\star$) or round (denoted as $\circ$) groups, reflecting the classifications used in the original kiki-bouba experiment.
@@ 3.1 Generating Images
To generate images, we first defined two sets of adjectives reflecting object appearance, $\mathcal{A}_{\circ}$ for roundness and $\mathcal{A}_{\star}$ for sharpness:
$\mathcal{A}_{\circ}$ =$=\{round,circular,soft,fat,chubby,$
@@
$\mathcal{A}_{\star}$ =$=\{sharp,spiky,angular,jagged,hard,$
@@
Then, we prompted a state-of the art text-to-image model [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1 using the template $\mathcal{P}$, where $\langle w\rangle\in\mathcal{A}_{\circ}\cup\mathcal{A}_{\star}$ is the adjective:
@@
For each adjective, we generated 25 images with different random seeds. This resulted in two collections of images, $\mathcal{I}_{\circ}$ and $\mathcal{I}_{\star}$, each comprising 250 images. Figure 2 displays several examples from these generated images.
Fig. 2: Examples of generated images: the upper row is from the sharp image set $\mathcal{I}_{\star}$, while the lower row is from the round image set $\mathcal{I}_{\circ}$.
@@ 3.2 Synthesizing Audios
Similar to the image generation process, we aim to collect two sets of audio samples, $\mathcal{S}_{\circ}$ for round sounds and $\mathcal{S}_{\star}$ for sharp sounds. We started with categorizing English consonants and vowels into round sounding or sharp sounding groups based on human preference alignments from prior work [52]:
@@
Finally, we utilized commercial text-to-speech models [^3] https://cloud.google.com/text-to-speech to generate audio. With each sequence synthesized with four distinct speaker identities, we yield two sets of audio samples, $\mathcal{S}_{\circ}$ and $\mathcal{S}_{\star}$, each containing 1944 samples.
@@ 4 Evaluation Method
To examine the presence of sound symbolism in the inherent knowledge of a pre-trained audio-visual model, we adopt a non-parametric approach to probe the model in a zero-shot setting. More specifically, we calculate a geometric score $s_{g}$ and a phonetic score $s_{p}$ using the synthesized dataset. [^4] Our evaluation framework is conceptually similar to that of [51]. However, it’s important to highlight that the scores in [51] are based solely on single-modality (text) inputs, focusing on the semantics implicitly exhibited in the surface forms of pseudowords. In contrast, our approach uses both audio and visual stimuli to probe the model, aiming to explore semantic similarity across multiple modalities. Our intuition is to identify a one-dimensional semantic direction within the model’s shared embedding space that best discriminates between sharp and round attributes. We then project query embeddings onto this direction to obtain a score representing their degree of association with each attribute. In each pre-trained model, there exists two modules, $\mathrm{F}_{a}\left(\cdot\right)$ and $\mathrm{F}_{i}\left(\cdot\right)$, which encode audio and images, respectively, into a shared embedding space.
Geometric Score: We start with extracting the embeddings of “round” and “sharp” images $\mathcal{I}_{\circ}$ and $\mathcal{I}_{\star}$ from the pre-trained model and identify the semantic direction in interest $\mathbf{w}_{g}$ in the embedding space:
$\mathbf{w}_{g}=\frac{1}{|\mathcal{I}_{\circ}|}\sum_{i\in\mathcal{I}_{\circ}}{\mathrm{F}_{i}\left(i\right)}-\frac{1}{|\mathcal{I}_{\star}|}\sum_{i\in\mathcal{I}_{\star}}{\mathrm{F}_{i}\left(i\right)}$ (1)
Then, for each sound $s\in\mathcal{S}_{\circ}\cup\mathcal{S}_{\star}$, we define the geometric score as the cosine similarity between its extracted embedding $\mathrm{F}_{a}\left(s\right)$ and $\mathbf{w}_{g}$.
@@
Phonetic Score: The phonetic score is calculated similarly to the geometric score, but using round and sharp sounds instead of images. Specifically, we feed ”round” and ”sharp” sounds ($\mathcal{S}_{\circ}$ and $\mathcal{S}_{\star}$) into our model to obtain two groups of embeddings and calculate the semantic direction:
$\mathbf{w}_{p}=\frac{1}{|\mathcal{S}_{\circ}|}\sum_{s\in\mathcal{S}_{\circ}}{\mathrm{F}_{a}\left(s\right)}-\frac{1}{|\mathcal{S}_{\star}|}\sum_{s\in\mathcal{S}_{\star}}{\mathrm{F}_{a}\left(s\right)}$ (3)
@@ 5.1 Pre-trained Audio-Visual Models
CAV-MAE [12] first extends single-modality MAE into a audio-visual multi-modality learner, followed by integrating contrastive learning and masked autoencoding into a single framework for enhanced representation. This approach simultaneously learns audio-visual pair information and audio-visual correspondence, facilitating the learning of joint and coordinated representations. AudioSet is used in pre-training.
@@
ImageBind [27] adopts an approach similar to CLIP, learning a joint embedding across six different modalities: images, text, audio, depth, thermal, and IMU data. It achieves this by binding them solely with image-paired data via contrastive loss. This method allows ImageBind to create a unified embedding space where inputs from diverse modalities are effectively aligned and correlated. The pre-training dataset is a combination of five datasets [54, 61, 60, 62, 63].
@@ 5.2 Quantitative Results
Fig. 3: Phones sorted by average geometric score grouped by the first syllable of the sounds. The colors indicates ground-truth association of each phone. (blue for $\circ$ and red for $\star$). Consonants and vowels are displayed on separate scales but are positioned absolutely to each other within each scale.
@@ 6 Conclusion
In this work, we investigate the presence of the kiki-bouba effect in pre-trained audio-visual models. Using a non-parametric approach in a zero-shot setting with specialized data, we probe the inherent knowledge of these models. Our results reveal a significant correlation between the models’ outputs and established patterns of sound symbolism in some audio-visual models. Specifically, audio-visual models trained on spoken image captions exhibit kiki-bouba effect patterns similar to those observed in humans. In contrast, models trained with general audio events show random or reversed tendencies. Additionally, using speech input demonstrates a more pronounced sound symbolism effect than using text forms. Our findings align with psychological literature, providing computational evidence for sound symbolism in audio-visual neural models.
@@ 7 Limitations
Additionally, our current analysis is limited to audio-visual models pre-trained on English-language corpora. In subsequent studies, we intend to evaluate models pre-trained on corpora from diverse languages, which could provide additional insights into the language-agnostic properties of sound symbolism, as suggested by previous studies [4, 6, 7].

### v3

@@ 1 Introduction
The field of deep learning has sustained significant interest in exploring the interplay between auditory and visual perceptions, two crucial modalities of human sensory processing. Known as audio-visual learning [1], this research field seeks to address the limitations of single-modality learning, enhancing performance on existing applications while opening new avenues for research. A key area of focus within audio-visual learning is representation learning, which aims to develop integrated audio-visual representations. Methods in this area often involve pretraining models to maximize the mutual information between auditory and visual inputs, whether in image or video form.
By training on diverse audio-visual datasets using specific algorithms, these models develop various specializations, such as speech-image retrieval [2, 3], visual speech recognition [4], and audio event understanding [5, 6]. These pre-trained models can extract highly informative, modality-agnostic representations from raw data, making them highly adaptable for audio-visual tasks, even with limited labeled data.
Beyond these specializations, some models have demonstrated intriguing language-processing capabilities. For instance, [4] shows that learning from speech with accompanying lip recordings enhances acoustic unit discovery. Additionally, [7, 8] demonstrates that audio-visual models trained solely on images paired with spoken captions, in a manner similar to human language acquisition, improves word discovery and segmentation abilities. Building on this, [9] further explored the potential of these models to perform syntactic parsing of spoken sentences.
@@
Building on these advancements, an intriguing research frontier lies in understanding how these models handle more abstract linguistic phenomena, such as sound symbolism [10, 11, 12, 13]. Sound symbolism refers to the systematic and non-arbitrary association between sounds and meanings—a phenomenon that challenges the traditional view of linguistic arbitrariness [14], suggesting that certain sounds naturally evoke specific meanings or visual qualities. For instance, the widely studied ”kiki-bouba effect,” illustrated in fig 1, demonstrates that people tend to associate the sharp-sounding word ”kiki” with angular shapes, while the softer-sounding ”bouba” is linked to rounder shapes [12, 13]. This phenomenon is robust across cultures and age groups [12, 15], indicating that some sound-meaning associations may be universally grounded in human perception.
According to the sound symbolism bootstrapping hypothesis for language acquisition [16], sound symbolism may play a foundational role in language learning by aiding toddlers in mapping and integrating multi-modal inputs, thereby supporting the formation of lexical representations. Since audio-visual models share certain parallels with human learning processes, we hypothesize that these models may naturally capture such associations. By investigating sound symbolism in these models, we can not only deepen our understanding of the cognitive parallels between human perception and artificial learning systems but also potentially enhance the models’ interpretability and alignment with human-like semantics in multimodal contexts. This exploration may thus open new directions for both linguistic theory and practical applications in audio-visual technology.
To explore this, we propose examining sound symbolism within audio-visual models by testing their ability to associate specific sounds with visual characteristics in ways consistent with human perception, as in the kiki-bouba effect. We start with curating a specialized dataset that includes synthesized images and audio samples, each distinctly categorized as ”sharp” or ”round.” Following [17], we then use a non-parametric approach in a zero-shot setting to probe the inherent knowledge of these models. Our experimental results reveal a significant correlation between the models’ outputs and established patterns of the kiki-bouba effect in some audio-visual models. Specifically, audio-visual models trained on spoken image captions are able to group visual stimuli and audio with different appearances significantly better than chance. Additionally, audio-visual models generally show a more pronounced sound symbolism effect than their purely textual vision-language model counterparts. Our findings show synergy with existing psychological literature [18, 19, 13, 20, 21] and further support the non-arbitrariness of human language.
@@ 2.3 Interpreting Deep Learning Models
Deep learning has undoubtedly achieved remarkable success across various fields [37, 38, 39]. Researchers are increasingly interested in interpreting these models to better understand the underlying mechanics and rationales behind their decisions [40]. This interpretation often involves feature visualization [41, 42, 43], saliency methods [44, 45, 43], and probing tasks [46, 47, 48, 49]. Recently, the natural language processing community has shown a growing interest in comparing the behavior of vision-language models to aspects of human language processing through probing tasks. These tasks assess models on compositionality [50], abstraction [51], and sound symbolism [17].
In this work, we specifically aim to study the effect of sound symbolism in audio-visual models. By integrating insights from the interpretability of deep learning models, we explore how these black-box systems may encode and reflect non-arbitrary associations between sounds and visual representations, akin to human perceptual and cognitive processes.
@@ 3 Dataset Collection
To assess the presence of sound symbolism in audio-visual models, we need to inspect how models behave with respect to sound and visual stimuli categorized into sharp or round groups. Therefore, we first construct a specialized dataset with images and audios grounded and grouped by human preference. While it is feasible to use controlled examples from original kiki-bouba experiments (as illustrated in Figure 1), the limited diversity in these samples could potentially undermine the reliability of our evaluation results. Hence, we enhanced the diversity by synthesizing data with generative models, as detailed in Sections 3.1 and 3.2. Overall, our dataset consists of 500 images and 3888 audio samples, each categorized into sharp or round groups, reflecting the classifications used in the original kiki-bouba experiment.
@@ 3.1 Generating Images
To generate images, we first defined two sets of adjectives reflecting object appearance, $\mathcal{W}_{\text{round}}$ for roundness and $\mathcal{W}_{\text{sharp}}$ for sharpness:
$\mathcal{W}_{\text{round}}$ =$=\{round,circular,soft,fat,chubby,$
@@
$\mathcal{W}_{\text{sharp}}$ =$=\{sharp,spiky,angular,jagged,hard,$
@@
Then, we prompted a state-of the art text-to-image model [^1] https://huggingface.co/stabilityai/stable-diffusion-2-1 using the template $\mathcal{P}$, where $\langle w\rangle\in\mathcal{W}_{\text{round}}\cup\mathcal{W}_{\text{sharp}}$ is the adjective:
@@
For each adjective, we generated 25 images with different random seeds. This resulted in two collections of images, $\mathcal{I}_{\text{round}}$ and $\mathcal{I}_{\text{sharp}}$, each comprising 250 images. Figure 2 displays several examples from these generated images.
Fig. 2: Examples of generated images: the upper row is from the sharp image set $\mathcal{I}_{\text{sharp}}$, while the lower row is from the round image set $\mathcal{I}_{\text{round}}$.
@@ 3.2 Synthesizing Audios
Similar to the image generation process, we aim to collect two sets of audio samples, $\mathcal{S}_{\text{round}}$ for round sounds and $\mathcal{S}_{\text{sharp}}$ for sharp sounds. We started with categorizing English consonants and vowels into round sounding or sharp sounding groups based on human preference alignments from prior work [52]:
@@
Finally, we utilized commercial text-to-speech models [^2] https://cloud.google.com/text-to-speech to generate audio. With each sequence synthesized with four distinct speaker identities, we yield two sets of audio samples, $\mathcal{S}_{\text{round}}$ and $\mathcal{S}_{\text{sharp}}$, each containing 1944 samples.
@@ 4 Evaluation Method
To examine the presence of sound symbolism in the inherent knowledge of a pre-trained audio-visual model, we adopt a non-parametric approach to probe the model in a zero-shot setting. More specifically, we calculate a geometric score $s_{g}$ and a phonetic score $s_{p}$ using the synthesized dataset, as similar to [17] [^3] Our evaluation framework is conceptually similar to that of [17]. However, it’s important to highlight that the scores in [17] are based solely on single-modality (text) inputs, focusing on the semantics implicitly exhibited in the surface forms of pseudowords. In contrast, our approach uses both audio and visual stimuli to probe the model, aiming to explore semantic similarity across multiple modalities.. Our intuition is to identify a one-dimensional semantic direction within the model’s shared embedding space that best discriminates between sharp and round attributes. We then project query embeddings onto this direction to obtain a score representing their degree of association with each attribute. In each pre-trained model, there exists two modules, $\mathrm{F}_{a}\left(\cdot\right)$ and $\mathrm{F}_{i}\left(\cdot\right)$, which encode audio and images, respectively, into a shared embedding space.
Geometric Score: We start with extracting the embeddings of “round” and “sharp” images $\mathcal{I}_{\text{round}}$ and $\mathcal{I}_{\text{sharp}}$ from the pre-trained model and identify the semantic direction in interest $\mathbf{w}_{g}$ in the embedding space:
$\mathbf{w}_{g}=\frac{1}{|\mathcal{I}_{\text{round}}|}\sum_{i\in\mathcal{I}_{\text{round}}}{\mathrm{F}_{i}\left(i\right)}-\frac{1}{|\mathcal{I}_{\text{sharp}}|}\sum_{i\in\mathcal{I}_{\text{sharp}}}{\mathrm{F}_{i}\left(i\right)}$ (1)
Then, for each sound $s\in\mathcal{S}_{\text{round}}\cup\mathcal{S}_{\text{sharp}}$, we define the geometric score as the cosine similarity between its extracted embedding $\mathrm{F}_{a}\left(s\right)$ and $\mathbf{w}_{g}$.
@@
Phonetic Score: The phonetic score is calculated similarly to the geometric score, but using round and sharp sounds instead of images. Specifically, we feed ”round” and ”sharp” sounds ($\mathcal{S}_{\text{round}}$ and $\mathcal{S}_{\text{sharp}}$) into our model to obtain two groups of embeddings and calculate the semantic direction:
$\mathbf{w}_{p}=\frac{1}{|\mathcal{S}_{\text{round}}|}\sum_{s\in\mathcal{S}_{\text{round}}}{\mathrm{F}_{a}\left(s\right)}-\frac{1}{|\mathcal{S}_{\text{sharp}}|}\sum_{s\in\mathcal{S}_{\text{sharp}}}{\mathrm{F}_{a}\left(s\right)}$ (3)
@@ 5.1 Pre-trained Audio-Visual Models
CAV-MAE [5] first extends single-modality MAE into a audio-visual multi-modality learner, followed by integrating contrastive learning and masked autoencoding into a single framework for enhanced representation. This approach simultaneously learns audio-visual pair information and audio-visual correspondence, facilitating the learning of joint and coordinated representations. AudioSet is used in pre-training. We select this model to test sound symbolism in audio-visual model that isn’t focusing on speech in pretraining.
@@
ImageBind [29] adopts an approach similar to CLIP, learning a joint embedding across six different modalities: images, text, audio, depth, thermal, and IMU data. It achieves this by binding them solely with image-paired data via contrastive loss. This method allows ImageBind to create a unified embedding space where inputs from diverse modalities are effectively aligned and correlated. The pre-training dataset is a combination of five datasets [54, 61, 60, 62, 63]. We include this model to test whether including multiple modalities together help model to improve sound symbolism.
@@ 5.2 Quantitative Results
Fig. 3: Phones sorted by average geometric score grouped by the first syllable of the sounds. The colors indicates ground-truth association of each phone. (blue and $\circ$ refer to round group, while red and $\star$ represent sharp). Consonants and vowels are displayed on separate scales but are positioned absolutely to each other within each scale.
@@ 6 Conclusion
In this work, we investigate the presence of the kiki-bouba effect in pre-trained audio-visual models. Using a non-parametric approach in a zero-shot setting with specialized data, we probe the inherent knowledge of these models. Our results reveal a significant correlation between the models’ outputs and established patterns of sound symbolism in some audio-visual models. Specifically, audio-visual models trained on spoken image captions exhibit kiki-bouba effect patterns similar to those observed in humans. In contrast, models trained with general audio events show random or reversed tendencies. Additionally, using speech input demonstrates a more pronounced sound symbolism effect than using text forms. Our findings align with prior psychological literature, contribute valuable insights regarding the non-arbitrariness of language in linguistic theory.
@@ 7 Limitations
Additionally, our current analysis is limited to audio-visual models pre-trained on English-language corpora. In subsequent studies, we intend to evaluate models pre-trained on corpora from diverse languages, which could provide additional insights into the language-agnostic properties of sound symbolism, as suggested by previous studies [12, 15].